A caterer made an ice cream punch by combining fruit juice that costs $3.00 per gallon with ice cream that costs $6.00 per gallon. How many gallons of each were used to make 150 gal of punch costing $3.60 per gallon?
Answer by jorel555(1290)

Let n be the amount of ice cream needed. Then:
6n+3(150-n)=3.60(150)
3n+450=540
3n=90
n=30
The caterer needs 30 gallons of ice cream, and 120 gallons of punch. ☺☺☺☺